11.2. Alton Inc. is working at full production capacity producing 28,000 units of a unique product. Manufacturing costs per unit for the product are as follows:

Direct materials $9
Direct labor 8
Manufacturing overhead 10
Total manufacturing cost per unit $27

The per-unit manufacturing overhead cost is based on a $4 variable cost per unit and $168,000 fixed costs. The nonmanufacturing costs, all variable, are $10 per unit, and the sales price is $42 per unit.

Sports Headquarters Company (SHC) has asked Alton to produce 6,600 units of a modification of the new product. This modification would require the same manufacturing processes. However, because of the nature of the proposed sale, the estimated nonmanufacturing costs per unit are only $5 (not $10). Alton would sell the modified product to SHC for $32 per unit.

Required

1-a. Calculate the contribution margin for 6,600 units for both the current and special order.

Current = $72,600

Special Order = ?

2. Suppose that Alton Inc. had been working at less than full capacity to produce 23,200 units of the product when SHC made the offer. What is the minimum price per unit that Alton should accept for the modified product under these conditions?
Answer #1

Answer 1 (a):

Contribution margin for 6,600 units for the current order = $72,600

Contribution margin for 6,600 units for the special order = $39,600

Working:

Current Special Order $32 Selling Price Variable costs: Direct material Direct labor Variable Manufacturing overhead Variable

Answer 2:

Full production capacity = 28,000 units

working at capacity = 23,200 Units

Hence idle capacity = 28000 - 23200 =4,800 units

Special order is for 6,600 units

If Alton accepts the SHC order, it will lose existing sales of = 6600 - 4800 = 1800 units

Contribution lost due to loss of existing sales = 1800 * 11 = $19,800

Per unit additional contribution required from SHC order to make it up = 19800 /6600 = $3

Minimum price per unit that Alton should accept for the modified product = Variable cost per unit of SHC order + Additional contribution required

= $26 + $3

= $29

Minimum price per unit that Alton should accept for the modified product = $29